Alpha manufactures wheels at their Windsor facility. For next month the production manager plans on producing 4,000 wheels per day. The company can produce as many as 7,000 wheels per day, but are more likely to produce 6,000 per day. The demand for wheels for the next three years is expected to average 5,000 per day. Fixed manufacturing costs per month total $360,000. The company works 23 days a month due to local zoning restrictions. Fixed manufacturing overhead is charged on a per wheel basis.

Required:

a. What is the theoretical fixed manufacturing overhead rate per wheel?

b. What is the practical fixed manufacturing overhead rate per wheel?

c. What is the normal fixed manufacturing overhead rate per wheel?

d. What is the master-budget fixed manufacturing overhead rate per wheel?
Answer #1

a.Theoretical overhead rate = $360000 / (7,000 × 23) = $2.236

b.Practical overhead rate = $360000 / (6000 × 23) = $2.609

c.Normal overhead rate = $360000 / (5000 × 23) = $3.130

d.Master-budget overhead rate = $360000 / (4,000 × 23) = $3.913